# Amortized Bayesian Workflow

### v2

@@ 2.2.3 Step 3: Many-chains MCMC with amortized initializations
If PSIS does not yield satisfactory samples, we resort to an MCMC sampling scheme as a safe fall-back option. In our amortized workflow, the MCMC step is augmented by reusing computations from the previous steps as initialization values. In principle, this step can incorporate any MCMC algorithm suited to the problem at hand. Examples include slice sampling for models with non-differentiable likelihoods [55], or HMC [56] samplers when gradients are available.
In this work, we use the ChEES-HMC algorithm [36] as an instantiation of MCMC. Most notably, ChEES-HMC supports the execution of thousands of parallel chains on a GPU for high-throughput sampling [81]. Amortized posterior draws from previous steps provide a natural and convenient choice for initializing MCMC chains to accelerate convergence (Figure 4). This approach is conceptually similar to using methods like parallel quasi-Newton variational inference (i.e., Pathfinder; [100]) to obtain initial values for MCMC chains. However, the amortized initial values are drawn in parallel in near-instant time, while Pathfinder requires re-fitting the variational approximation for each new observed dataset. For the purpose of ChEES-HMC initializations with multimodal posterior distributions, it is again desirable that the amortized posterior draws are typically mass-covering (cf. Step 2). See the appendix for additional details on the ChEES-HMC algorithm.
@@
In this last step, we use the nested $\widehat{R}$ diagnostic [50], which is specifically designed to assess the convergence of the many-but-short MCMC chains. If the diagnostics in this step indicate unreliable inference, we recommend resorting to the overarching Bayesian workflow [27] and addressing the computational issues that even persist when using the (ChEES-)HMC algorithm. This could involve increasing the number of warmup iterations, using the established NUTS-HMC algorithm [37, 8], or revising the Bayesian model specification and parametrization.
Step 3: If (nested) $\widehat{R}$ is below the convergence threshold (e.g., $1.01$), accept the MCMC draws. Otherwise, increase warm-up or revise the model according to the standard Bayesian workflow.
@@ 2.3 Related work
Both simulation-based inference and amortized inference have seen rapid progress over the past decade [99, 12, 41], driven by the need to perform Bayesian inference in complex models with intractable likelihoods [e.g., 17, 92, 101, 29, 89, 3, 63]. These advances have been fueled by modern generative modeling, such as normalizing flows [60, 62, 32], transformers [54, 9, 93], diffusion models [79, 77, 45, 25, 30], consistency models [80, 73], and flow matching [46, 95].
To address the potential systematic errors of (amortized) neural posteriors, several works propose corrections using importance reweighting schemes [13, 82], augmented training objectives [14, 52, 58, 72], or post-hoc corrections [78]. Simultaneously, hybrid approaches that combine density estimators with MCMC have gained traction [70, 35, 24, 51, 2, 6, 33]. These include using variational approximations or learned flows as preconditioners for MCMC [35, 5], adaptive proposal mechanisms [61, 24], and initialization strategies to accelerate convergence or improve diagnostics [100, 90, 82].
Our proposed workflow builds on these lines of work by integrating amortized inference, likelihood-based correction, and many-chain MCMC into a unified, diagnostic-driven pipeline. Unlike traditional SBI pipelines, our workflow dynamically adapts the inference strategy to the dataset at hand, improving robustness without retraining. This modular design provides a practical foundation for principled amortized inference across diverse data regimes.
@@ 3.1 Procedure
For each problem, we begin by training the amortized posterior approximator on simulated parameter-observation pairs (i.e., simulation-based training). We verify that the model performance is satisfactory in a closed-world setting, as diagnosed by simulation-based calibration and parameter recovery checking (see Section 2.1). Details on diagnostic results, simulation budgets, and training hyperparameters are provided in the appendix.
@@
For the out-of-distribution diagnostics in Step 1, we use the $1-\alpha=95\%$ percentile as the rejection threshold. We compute Mahalanobis distances in the summary statistics using 10,000 training simulations. We draw 2,000 posterior samples from the amortized posterior $q_{\phi}$ at Step 1. In Step 2, we correct the amortized draws using PSIS, rejecting draws if Pareto-$\hat{k}>0.7$. Step 3 uses ChEES-HMC with convergence determined by nested $\widehat{R}<1.01$. We run 2048 chains in parallel (16 superchains, each with 128 subchains), with 200 warmup steps and a single sampling step, for a total of 2048 posterior draws.
@@
To assess the quality of posterior draws from our workflow, we compare them to reference posterior draws using two evaluation metrics: the 1-Wasserstein distance (W1) and the mean marginal total variation distance (MMTV). The W1 distance quantifies the overall discrepancy between full joint distributions. MMTV measures the lack of overlap between marginal distributions and takes value in the range $[0,1]$; for example, an MMTV value of 0.2 implies that, on average, the approximate posterior draws and reference draws share an 80% overlap for their marginal distributions. For both metrics, lower values indicate better posterior approximation quality. As a rule of thumb, MMTV values below 0.2 indicate good posterior approximation fidelity [1, 43].
@@ 3.2 Applications
We apply the proposed workflow to four posterior inference problems, including both simulated benchmarks and real-world experimental datasets. These case studies were chosen to reflect a range of commonly encountered statistical inference scenarios, including classical distributional parameter estimation and analyses of large-scale datasets arising in psychology and cognitive modeling. We describe each problem briefly below, with further details provided in the appendix.
@@
where $\mu\in\mathbb{R}$ is the location, $\sigma\in\mathbb{R}_{>0}$ is the scale, and $\xi\in\mathbb{R}$ is the shape parameter. We follow the prior specification from Caprani [7]. For each dataset, we collect $N=65$ i.i.d. observations and infer the posterior distribution over the parameter vector $\theta=(\mu,\sigma,\xi)$. We generate a total of $K=1000$ test datasets by deliberately simulating from a model with a 2$\times$ wider prior distribution to emulate out-of-distribution settings in real applications (see the appendix for details).
@@
The Bernoulli generalized linear model (GLM) is a classical model with binary outcomes, included in the SBI benchmark suite [48]. Each observation $y_{i}\in\{0,1\}$ is modeled as:
@@
where $v_{i}\in\mathbb{R}^{10}$ is a fixed input vector, $\theta\in\mathbb{R}^{10}$ is the parameter vector, and $\sigma(\cdot)$ denotes the logistic function. We generate $K=10,000$ in-distribution test datasets by sampling parameters from the model prior and simulating corresponding observations $\{y_{i}\}_{i=1}^{100}$ [48].
@@
Psychometric functions are widely used in perceptual and cognitive science to characterize the relationship between stimulus intensity and the probability of a specific response [94]. We use the overdispersed hierarchical model from Schütt et al. [75], where the number of correct trials $y_{i}$ at stimuli level $x_{i}$ is modeled as:
@@
where $n_{i}$ is the number of trials, $\eta\in[0,1]$ controls overdispersion, and $\bar{p}_{i}=\psi(x_{i};m,w,\lambda,\gamma)$ is the expected success probability given by the psychometric function $\psi(x;m,w,\lambda,\gamma)=\gamma+(1-\lambda-\gamma)\,S(x;m,w)$, where $S$ is a sigmoid function (e.g., cumulative normal), $m$ is the threshold, $w$ is the width, $\lambda$ is the lapse rate for infinitely high stimulus levels, and $\gamma$ is the guess rate for infinitely low stimulus levels. In total, the model parameters are $\theta=(m,w,\lambda,\gamma,\eta)$. Our empirical evaluation uses 8,526 mouse behavioral datasets from the International Brain Laboratory public database [85].
@@
The drift-diffusion model (DDM) is a popular evidence accumulation model for psychological models of human decision making [65]. It describes a two-choice decision task as a stochastic process in which noisy evidence accumulates over time until it reaches one of the decision boundaries. The evolution of the decision variable $z(t)$ is modeled as
@@
where $v$ is the drift rate (the average rate of evidence accumulation), $\sigma$ is the noise scale, and $W(t)$ denotes a standard Wiener process. A decision is made when $z(t)$ reaches either a positive or negative boundary, typically placed symmetrically at ±$\pm a$, where $a$ is the boundary separation. The model also includes a non-decision time parameter $\tau$, capturing processes that are not part of the decision process. We adopt the model specification from von Krause et al. [89], which extends the standard DDM to incorporate experimental condition effects via six parameters: $\theta=(v_{1},v_{2},a_{1},a_{2},\tau_{c},\tau_{n}$). The test datasets consist of 15,000 participants from the online implicit association test (IAT) database [96, 89], providing a large-scale, real-world benchmark for Bayesian inference in cognitive modeling.
@@ 3.3 Main Results
Table 1: Summary of our amortized Bayesian workflow across four problems. For each step, we report the number of accepted datasets, wall-clock time (minutes), and time per accepted dataset (TPA) in seconds. The time for Step 1 includes training, inference, and diagnostics for the amortized model. “Workflow total” aggregates the results of our method across all steps. As a baseline reference, “Baseline workflow total” is an estimate of the total required runtime for ChEES-HMC on all datasets.
@@
GEV Step 1: Amortized inference $523/1000$ 3 0.4
@@
Bernoulli GLM Step 1: Amortized inference $9519/10000$ 1 0.007
@@
Step 3: ChEES-HMC w/ inits $/56$ 56 4 4
Workflow total (ours) $/10000$ 10000 5 0.03
@@
Psychometric curve Step 1: Amortized inference $7213/8526$ 7 0.06
@@
Decision model Step 1: Amortized inference $13498/15000$ 86 0.4
@@
Figure 5: Evaluation of posterior draws across four problems based on two metrics: W1 distance (top row) and MMTV distance (bottom row). Lower values indicate better posterior approximation. ABI(✓) and ABI(✗) denote accepted and rejected draws, respectively, from amortized Bayesian inference in Step 1. PSIS denotes importance-weighted draws accepted in Step 2, and C-HMC denotes draws accepted via ChEES-HMC in Step 3.
Table 1 summarizes the performance of the proposed amortized Bayesian workflow across the four problems described in Section 3.2. Step 1 (ABI) exhibits extremely low time per accepted dataset (TPA), with most of the cost incurred as a one-time expense during the training phase—including prior simulation, model training, and diagnostic evaluation. Once trained, ABI incurs negligible marginal cost ($\ll 1\text{sec}$) when applied to a new dataset. Datasets flagged as out-of-distribution in Step 1 are forwarded to Step 2 for correction via PSIS. PSIS is highly effective, successfully correcting most rejected amortized draws and substantially reducing the number of datasets requiring full MCMC. Only a small subset of datasets progresses to Step 3, where ChEES-HMC is used for high-fidelity sampling. As the most computationally expensive component, ChEES-HMC is applied selectively, allowing the workflow to retain both accuracy and efficiency. Overall, the amortized workflow completes inference for nearly all datasets. [^2] A small number of datasets with particularly difficult properties require extended MCMC runs to converge. Compared to using ChEES-HMC for all datasets, our workflow achieves substantial computational savings—approximately over 5×5$\times$5 ×, 120×120$\times$120 ×, 60×60$\times$60 ×, and 15×15$\times$15 × faster for the GEV, Bernoulli GLM, psychometric curve, and decision model tasks, respectively.
@@ 3.4 Advantage of amortized initializations for MCMC
To evaluate whether amortized posterior estimates remain useful in such cases, we test their effectiveness as initializations for ChEES-HMC chains. We conduct experiments on 20 randomly selected test datasets that progress to Step 3 of the workflow. This indicates that both the amortized posterior draws and their Pareto-smoothed refinement are deemed unacceptable, as quantified by Pareto-$\hat{k}>0.7$ in Step 2. We compare three initialization methods for ChEES-HMC chains: (1) amortized posterior draws, (2) PSIS-refined amortized draws, and (3) a random initialization scheme similar to Stan [8]. We run the chains for varying numbers of warmup iterations, followed by a single sampling iteration. As described in Section 2, we use the nested $\widehat{R}$ value to gauge whether the chains converged appropriately during the warmup stage, as quantified by the common $\widehat{R}-1$ threshold of $0.01$ [87].
Figure 6: Using amortized posterior draws as initializations for ChEES-HMC reduces the required warmup in the GEV and decision model tasks. We show median±$\pm$±IQR across 20 test datasets in Step 3.
@@
Overall, these results demonstrate that amortized inference may provide suitable initializations for ChEES-HMC. However, the added benefit of initializing chains with PSIS-refined amortized draws (Step 2) instead of raw amortized draws (Step 1) remains unclear. While PSIS often accelerates convergence, it occasionally degrades worst-case performance (see upper error bounds for GEV task in Figure 6). We further study the impact of initialization for the popular NUTS sampler [37], with similar results: amortized initializations reduce the required warmup in most cases (see appendix).

### v3

@@ 2.2.3 Step 3: Many-chains MCMC with amortized initializations
If PSIS does not yield satisfactory results, we resort to an MCMC sampling scheme as a safe fallback option. In our amortized workflow, the MCMC step is augmented by reusing computations from the previous steps as initialization values. In principle, this step can incorporate any MCMC algorithm suited to the problem at hand. Examples include slice sampling for models with non-differentiable likelihoods (Neal, 2003), or HMC (Neal, 2011) samplers when gradients are available.
In this work, we use the ChEES-HMC algorithm (Hoffman et al., 2021) as an instantiation of MCMC. Most notably, ChEES-HMC supports the execution of thousands of parallel chains on a GPU for high-throughput sampling (Sountsov et al., 2024). Amortized posterior draws from previous steps provide a natural and convenient choice for initializing MCMC chains to accelerate convergence (Figure 4). This approach is conceptually similar to using methods like parallel quasi-Newton variational inference (i.e., Pathfinder; Zhang et al., 2022) to obtain initial values for MCMC chains. However, the amortized initial values are drawn in parallel in near-instant time, while Pathfinder requires re-fitting the variational approximation for each new observed dataset. For the purpose of ChEES-HMC initializations with multimodal posterior distributions, it is again desirable that the amortized posterior draws are typically mass-covering (cf. Step 2). See Appendix A for additional details on the ChEES-HMC algorithm.
@@
In this last step, we use the nested $\widehat{R}$ diagnostic (Margossian et al., 2024), which is specifically designed to assess the convergence of the many-but-short MCMC chains. [^5] In more conventional settings involving long MCMC chains, the standard $\widehat{R}$ diagnostics (Gelman & Rubin, 1992; Vehtari et al., 2021) can be applied. If the diagnostics in this step indicate unreliable inference, we recommend resorting to the overarching Bayesian workflow (Gelman et al., 2020) and addressing the computational issues that even persist when using the (ChEES-)HMC algorithm. This could involve increasing the number of warmup iterations, using the established NUTS-HMC algorithm (Hoffman & Gelman, 2014; Carpenter et al., 2017), or revising the Bayesian model specification and parametrization.
Step 3: If (nested) $\widehat{R}$ is below the convergence threshold (e.g., $1.01$), accept the MCMC draws. Otherwise, increase warm-up or revise the model according to the standard Bayesian workflow (Gelman et al., 2020).
@@ 2.3 Related work
Both simulation-based inference and amortized inference have seen rapid progress over the past decade (Zammit-Mangion et al., 2025; Cranmer et al., 2020; Lavin et al., 2021), driven by the need to perform Bayesian inference in complex models with intractable likelihoods (e.g., Dingeldein et al., 2024; Wehenkel et al., 2024; Zhou et al., 2024; Ghaderi-Kangavari et al., 2023; von Krause et al., 2022; Bieringer et al., 2021; Radev et al., 2021). These advances have been fueled by modern generative modeling, such as normalizing flows (Papamakarios et al., 2021; Radev et al., 2020; Greenberg et al., 2019), transformers (Müller et al., 2022; Chang et al., 2025; Whittle et al., 2025), diffusion models (Song et al., 2021; Sharrock et al., 2024; Linhart et al., 2024; Geffner et al., 2023; Gloeckler et al., 2024), consistency models (Song et al., 2023; Schmitt et al., 2024b), and flow matching (Lipman et al., 2023; Wildberger et al., 2023). Practical software toolkits such as BayesFlow (Radev et al., 2023) and sbi (Boelts et al., 2025) further make these simulation-based inference techniques accessible to practitioners in user-friendly interfaces.
To address the potential systematic errors of (amortized) neural posteriors, several works propose corrections using importance reweighting schemes (Dax et al., 2023; Starostin et al., 2025), augmented training objectives (Delaunoy et al., 2022; Mishra et al., 2025; Orozco et al., 2025; Schmitt et al., 2024a), or post-hoc corrections (Siahkoohi et al., 2023). Simultaneously, hybrid approaches that combine density estimators with MCMC have gained traction (Salimans et al., 2015; Hoffman et al., 2019; Gabrié et al., 2022; Midgley et al., 2022; Arbel et al., 2021; Cabezas et al., 2024; Grenioux et al., 2023). These include using variational approximations or learned flows as preconditioners for MCMC (Hoffman et al., 2019; Cabezas & Nemeth, 2023), adaptive proposal mechanisms (Parno & Marzouk, 2018; Gabrié et al., 2022), and initialization strategies to accelerate convergence or improve diagnostics (Zhang et al., 2022; Wang et al., 2023; Starostin et al., 2025).
More broadly, automated Bayesian inference has been a central design goal of probabilistic programming systems such as Stan (Carpenter et al., 2017), PyMC (Oriol et al., 2023), (Num)Pyro (Bingham et al., 2019; Phan et al., 2019). These libraries provide general-purpose inference engines—typically gradient-based MCMC and variational inference—that can be applied to a wide range of likelihood-based models and are accompanied by well-developed diagnostic recommendations and workflow guidelines (Gelman et al., 2020). However, they do not natively support amortized inference across many datasets, and inference must be rerun from scratch for each dataset.
Our proposed workflow builds on and complements these lines of work by integrating amortized inference, likelihood-based correction, and many-chain MCMC into a unified, modular, and diagnostic-driven pipeline for accelerating Bayesian inference. It dynamically adapts the inference strategy to the dataset at hand—using amortized posterior draws when they are adequate and escalating to PSIS and MCMC otherwise—thereby improving the robustness of amortized inference and the overall efficiency of posterior computation. This modular design provides a practical foundation for principled amortized inference across diverse data regimes.
@@ 3.1 Procedure
For each problem, we begin by training the amortized posterior estimator on simulated parameter-observation pairs (i.e., simulation-based training). We verify that the model performance is satisfactory in a closed-world setting, as diagnosed by simulation-based calibration and parameter recovery checking (see Section 2.1). Details on diagnostic results, simulation budgets, and training hyperparameters are provided in Appendix C.
@@
For the out-of-distribution diagnostics in Step 1, we use the $\alpha=0.05$ as the rejection threshold. We compute Mahalanobis distances in the summary statistics using 10,000 training simulations. We draw 2,000 posterior samples from the amortized posterior $q_{\phi}$ at Step 1. In Step 2, we correct the amortized draws using PSIS, rejecting draws if Pareto-$\hat{k}>0.7$. Step 3 uses ChEES-HMC with convergence determined by nested $\widehat{R}<1.01$. We run 2048 chains in parallel (16 superchains, each with 128 subchains), with 200 warmup steps and a single sampling step, for a total of 2048 posterior draws.
@@
To assess the quality of posterior draws from our workflow, we compare them to reference posterior draws using two evaluation metrics: the 1-Wasserstein distance (W1) and the mean marginal total variation distance (MMTV). The W1 distance quantifies the overall discrepancy between full joint distributions. MMTV measures the lack of overlap between marginal distributions and takes value in the range $[0,1]$; for example, an MMTV value of 0.2 implies that, on average, the approximate posterior draws and reference draws share an 80% overlap for their marginal distributions. For both metrics, lower values indicate better posterior approximation quality. As a rule of thumb, MMTV values below 0.2 indicate good posterior approximation fidelity (Acerbi, 2020; Li et al., 2025).
@@ 3.2 Applications
We apply the proposed workflow to four posterior inference problems, including both simulated benchmarks and real-world experimental datasets. These case studies were chosen to reflect a range of commonly encountered statistical inference scenarios, including classical distributional parameter estimation and analyses of large-scale datasets arising in psychology and cognitive modeling. We describe each problem briefly below, with further details provided in Appendix C.
@@
[…]
@@
The Bernoulli generalized linear model (GLM) is a classical model with binary outcomes, included in the SBI benchmark suite (Lueckmann et al., 2021). Each observation $y_{i}\in\{0,1\}$ is modeled as:
@@
where $v_{i}\in\mathbb{R}^{10}$ is a fixed input vector, $\theta\in\mathbb{R}^{10}$ is the parameter vector, and $\sigma(\cdot)$ denotes the logistic function. We generate $K=10,000$ in-distribution test datasets by sampling parameters from the model prior and simulating corresponding observations $\{y_{i}\}_{i=1}^{100}$ (Lueckmann et al., 2021).
@@
Psychometric functions are widely used in perceptual and cognitive science to characterize the relationship between stimulus intensity and the probability of a specific response (Wichmann & Hill, 2001). We use the overdispersed hierarchical model from Schütt et al. (2016), where the number of correct trials $y_{i}$ at stimuli level $x_{i}$ is modeled as:
@@
where $n_{i}$ is the number of trials, $\eta\in[0,1]$ controls overdispersion, and $\bar{p}_{i}=\psi(x_{i};m,w,\lambda,\gamma)$ is the expected success probability given by the psychometric function $\psi(x;m,w,\lambda,\gamma)=\gamma+(1-\lambda-\gamma)\,S(x;m,w)$, where $S$ is a sigmoid function (e.g., cumulative normal), $m$ is the threshold, $w$ is the width, $\lambda$ is the lapse rate for infinitely high stimulus levels, and $\gamma$ is the guess rate for infinitely low stimulus levels. In total, the model parameters are $\theta=(m,w,\lambda,\gamma,\eta)$. Our empirical evaluation uses 8,526 mouse behavioral datasets from the International Brain Laboratory public database (The International Brain Laboratory et al., 2021).
@@
The drift-diffusion model (DDM) is a popular evidence accumulation model for psychological models of human decision making (Ratcliff & McKoon, 2008). It describes a two-choice decision task as a stochastic process in which noisy evidence accumulates over time until it reaches one of the decision boundaries. The evolution of the decision variable $z(t)$ is modeled as
@@
where $v$ is the drift rate (the average rate of evidence accumulation), $\sigma$ is the noise scale, and $W(t)$ denotes a standard Wiener process. A decision is made when $z(t)$ reaches either a positive or negative boundary, typically placed symmetrically at $\pm a$, where $a$ is the boundary separation. The model also includes a non-decision time parameter $\tau$, capturing processes that are not part of the decision process. We adopt the model specification from von Krause et al. (2022), which extends the standard DDM to incorporate experimental condition effects via six parameters: $\theta=(v_{1},v_{2},a_{1},a_{2},\tau_{c},\tau_{n}$). The test datasets consist of 15,000 participants from the online implicit association test (IAT) database (Xu et al., 2014; von Krause et al., 2022), providing a large-scale, real-world benchmark for Bayesian inference in cognitive modeling.
@@ 3.3 Main Results
Table 1: Summary of our amortized Bayesian workflow across four problems. For each step, we report the number of accepted datasets, wall-clock time (minutes), and time per accepted dataset (TPA) in seconds. The time for the training phase includes amortized estimator training, simulations, and diagnostics evaluations. The time for Step 1 includes amortized posterior draws and the OOD test. The TPA for Step 1 accounts for both the training phase and Step 1. “Workflow total” aggregates the results of our method across all steps. As a baseline reference, “Baseline workflow total” is an estimate of the total required runtime for ChEES-HMC on all datasets.
@@
GEV Training phase 3 0.4
Step 1: Amortized inference $523/1000$ 0.1
@@
Bernoulli GLM Training phase 0.8 0.007
Step 1: Amortized inference $9519/10000$ 0.3
@@
Step 3: ChEES-HMC w/ inits $56/56$ 4 4
Workflow total (ours) $10000/10000$ 5 0.03
@@
Psychometric curve Training phase 6 0.06
Step 1: Amortized inference $7213/8526$ 0.4
@@
Decision model Training phase 85 0.4
Step 1: Amortized inference $13498/15000$ 1
@@
Figure 5: Evaluation of posterior draws across four problems based on two metrics: W1 distance (top row) and MMTV distance (bottom row). Lower values indicate better posterior approximation. ABI(✓) and ABI(✗) denote accepted and rejected draws, respectively, from amortized Bayesian inference in Step 1. PSIS denotes importance-weighted draws accepted in Step 2, and C-HMC denotes draws accepted via ChEES-HMC in Step 3. Metrics are computed on up to 100 datasets for each type of draws.
Table 1 summarizes the performance of the proposed amortized Bayesian workflow across the four problems described in Section 3.2. Step 1 (ABI) exhibits extremely low time per accepted dataset (TPA), with most of the cost incurred as a one-time expense during the training phase—including prior simulation, model training, and diagnostic evaluation. Once trained, ABI incurs negligible marginal cost ($\ll 1\text{sec}$) when applied to a new dataset. Datasets flagged as out-of-distribution in Step 1 are forwarded to Step 2 for correction via PSIS. PSIS is highly effective, successfully correcting most rejected amortized draws and substantially reducing the number of datasets requiring full MCMC. Only a small subset of datasets progresses to Step 3, where ChEES-HMC is used for high-fidelity sampling. As the most computationally expensive component, ChEES-HMC is applied selectively, allowing the workflow to retain both accuracy and efficiency. Overall, the amortized workflow completes inference for nearly all datasets. [^6] A small number of datasets with particularly difficult properties require extended MCMC runs to converge. Compared to using ChEES-HMC for all datasets, our workflow achieves substantial computational savings—approximately over $5\times$, $120\times$, $60\times$, and $15\times$ faster for the GEV, Bernoulli GLM, psychometric curve, and decision model tasks, respectively.
@@ 3.4 Advantage of amortized initializations for MCMC
To evaluate whether amortized posterior estimates remain useful in such cases, we test their effectiveness as initializations for ChEES-HMC chains. We conduct experiments on 20 randomly selected test datasets that progress to Step 3 of the workflow. This indicates that both the amortized posterior draws and their Pareto-smoothed refinement are deemed unacceptable, as quantified by Pareto-$\hat{k}>0.7$ in Step 2. We compare three initialization methods for ChEES-HMC chains: (1) amortized posterior draws, (2) PSIS-refined amortized draws, and (3) a random initialization scheme similar to Stan (Carpenter et al., 2017). We run the chains for varying numbers of warmup iterations, followed by a single sampling iteration. As described in Section 2, we use the nested $\widehat{R}$ value to gauge whether the chains converged appropriately during the warmup stage, as quantified by the common $\widehat{R}-1$ threshold of $0.01$ (Vehtari et al., 2021).
[…]
@@
Overall, these results demonstrate that amortized inference may provide suitable initializations for ChEES-HMC. However, the added benefit of initializing chains with PSIS-refined amortized draws (Step 2) instead of raw amortized draws (Step 1) remains unclear. While PSIS often accelerates convergence, it occasionally degrades worst-case performance (see upper error bounds for GEV task in Figure 6). We further study the impact of initialization for the popular NUTS sampler (Hoffman & Gelman, 2014), with similar results: amortized initializations reduce the required warmup in most cases (see Appendix E).